### v2

# arXiv:2601.11839v2[math.FA] 28 Feb 2026

Abstract
For a crystal group $\Gamma$ in dimension $n$, a closed subspace $\mathcal{V}$ of $L^{2}(\mathbb{R}^{n})$ is called $\Gamma$–shift-invariant if, for every $f\in\mathcal{V}$, the shifts of $f$ by every element of $\Gamma$ also belong to $\mathcal{V}$. The main purpose of this paper is to provide a characterization of the $\Gamma$–shift-invariant closed subspaces of $L^{2}(\mathbb{R}^{n})$ that is valid for all crystal groups, non-symmorphic as well as symmorphic.
1 Introduction
Let $\Gamma$ be a crystal group in dimension $n$ with point group $\Pi$ and associated lattice $\mathcal{L}$. Detailed definitions are given in Section 3. Elements of $\Gamma$ are written $[x,M]$, where $M$ is an orthogonal matrix that belongs to $\Pi$ and $x$ is a vector in $\mathbb{R}^{n}$. The crystal group is called symmorphic if $0,M]\in\Gamma$, for all $M\in\Pi$; or, equivalently, if $\Gamma$ is isomorphic to the semi-direct product $\mathcal{L}\rtimes\Pi$. If this is not the case, then $\Gamma$ is non-symmorphic.
For a function $f$ on $\mathbb{R}^{n}$, the shift of $f$ by $[x,M]$ is $\pi[x,M]f$ given by
$\pi[x,M]f(z)=f\left(M^{-1}z-x\right),$
for $z\in\mathbb{R}^{n}$. When these shifts are applied to functions in the Hilbert space of all square-integrable functions, $L^{2}(\mathbb{R}^{n})$, they constitute a unitary representation $\pi$ of $\Gamma$, which we call the natural representation. A closed subspace $\mathcal{V}$ of $L^{2}(\mathbb{R}^{n})$ is called $\pi$-invariant, or $\Gamma$–shift-invariant, if $\pi[x,M]f\in\mathcal{V}$, for every $f\in\mathcal{V}$ and every $x,M]\in\Gamma$. Our purpose is to provide a characterization of the $\Gamma$–shift-invariant closed subspaces of $L^{2}(\mathbb{R}^{n})$.
The term shift-invariant closed subspace means a closed subspace $\mathcal{V}$ of $L^{2}(\mathbb{R}^{n})$ such that $T_{k}f\in\mathcal{V}$, for all $f\in\mathcal{V}$ and all $k\in\mathbb{Z}^{n}$, where $T_{k}f(z)=f(z-k)$, for $z\in\mathbb{R}^{n}$. In every dimension, there is one crystal group that is abelian and it is the set of shifts by lattice points; so shift-invariance is $\Gamma$–shift invariance when $\Gamma$ is isomorphic to $\mathbb{Z}^{n}$. Understanding shift-invariant subspaces became more important with the rise of wavelet analysis (see [10] for a general introduction to the classical theory of wavelets) as the central subspace in a multiresolution analysis is shift-invariant; see also [21] for more on the role of shift-invariant subspaces in wavelet theory. A characterization of shift-invariant subspaces was known in 1964 by Helson [15] and was reformulated in 2000 by Bownik [6] where he applied it to wavelet theory.
The theory of wavelets has many generalizations and variations. One generalization, wavelets with composite dilations, was introduced in [13] and [14]. The term composite dilations refers to the addition of another set $\mathcal{B}$ of matrices that are moving functions, besides the shifts by vectors in a lattice and the powers of a single dilating matrix. Often, $\mathcal{B}$ is a finite group of measure preserving matrices. Blanchard developed the theory in this case in [4]; see also [5] where the focus was on developing Haar-type wavelet systems when $\mathcal{B}$ leaves the shift lattice invariant and the semi-direct product of $\mathcal{B}$ with the lattice is actually a crystal group, necessarily symmorphic. In [17] MacArthur and one of the current authors introduced the concept of a multiresolution analysis where the shifts came from an arbitrary crystal group and cast the theory in the context of the abstract approach of [2]. Independently, González and Moure [12] also formulated the theory for shifts by a crystal group. In the definition of a multiresolution analysis or a generalized multiresolution analysis for either composite dilations, with a finite $\mathcal{B}$, or shifts by a crystal group, the central subspace in the multiresolution analysis is left invariant under shifts by a crystal group $\Gamma$.
Another reason to be concerned about the nature of $\Gamma$–shift-invariant closed subspaces is the emergence of topological quantum chemistry (see [9] and [8], for example) where band representations play a significant role. A band representation of a crystal group $\Gamma$ is the restriction of the natural representation $\pi$ to a closed $\Gamma$–shift-invariant subspace generated by a single function in $L^{2}(\mathbb{R}^{n})$. A readable review of band representations can be found in [1].
In 2020, Barbieri, et al, [3] considered $\Gamma$–shift-invariance of closed subspaces of $L^{2}(\mathcal{G})$, where $\mathcal{G}$ is a locally compact abelian group and $\Gamma$ is the semi-direct product, $\Lambda\rtimes G$, where $\Lambda$ is a uniform lattice in $\mathcal{G}$ and $G$ is a finite group of continuous automorphisms of $\mathcal{G}$ that fix $\Lambda$. Theorem 4.7 of [3] provides a characterization of $\Gamma$–shift-invariant closed subspaces of $L^{2}(\mathcal{G})$. When $\mathcal{G}=\mathbb{R}^{n}$, this generalizes the characterization of shift-invariant closed subspaces of $L^{2}(\mathbb{R}^{n})$ given in [6] to $\Gamma$–shift-invariance, where $\Gamma$ is a symmorphic crystal group.
Our main result, given in Theorem 6.3 and reformulated in Corollary 6.7, applies to all crystal groups. Thus, the contribution of this paper is to describe the situation for non-symmorphic crystal groups. The two crystal groups in one dimension are symmorphic, while 4 of the 17 two dimensional crystal groups are non-symmorphic. In three dimensions, 146 out of the 219 different (up to isomorphism) crystal groups are non-symmorphic. It gets more dramatic as the dimension increases. In dimensions 5 and 6, more than 85% and 97%, respectively, of the crystal groups are non-symmorphic (see [19] for exact numbers). A readable account of the process of enumeration of crystal groups is given in [18]. Also, Section 2.5 of [18] contains an explanation of the connection between a vector system for a given point group and the non-symmorphic crystal groups with that point group. We will discuss how a vector system can be defined for a particular $\Gamma$ in Section 3 and it turns out that the vector system plays a key role in our main result.
Imprecisely stated, we show, in Theorem 6.3 and Corollary 6.7, that there is a one-to-one correspondence between $\Gamma$–shift-invariant closed subspaces of $L^{2}(\mathbb{R}^{n})$ and certain maps (twisted range functions, where the twist comes from the vector system) from an open subset of $\mathbb{R}^{n}$ and the closed subspaces of $\ell^{2}(\mathcal{L}^{*})$, where $\mathcal{L}^{*}$ is the dual lattice of $\mathcal{L}$. If $\Gamma=\mathbb{Z}^{n}$, then the characterization in Theorem 6.3 reduces to the characterization of shift-invariant subspaces given in Proposition 1.5 in [6]. When $\Gamma$ happens to be symmorphic, Theorem 6.3 is equivalent to the characterization for symmorphic crystal groups found in [3]. Our approach draws on the presentation of range functions by Bownik and Ross in [7], to the extent that $\Gamma$–shift-invariance implies invariance under shifts by the lattice $\mathcal{L}$. We view $\pi$ as a unitary representation of $\Gamma$ and our strategy is to use unitary maps to transform $\pi$ into an equivalent representation that can be analyzed using range function methods.
We gather together preliminary facts and known results in Section 2 while Section 3 is devoted to organizing the properties of crystal groups that are needed. The unitary maps that will be used to transform $\pi$ are introduced in Section 4 and applied in the next section. In the final section, we prove our main theorem, Theorem 6.3. We also present an example of a non-symmorphic wallpaper group $\Gamma$ and a $\Gamma$–shift-invariant closed subspace of $L^{2}(\mathbb{R}^{2})$ where the impact of the non-symmorphic nature of this $\Gamma$ is illustrated and the twist coming from the vector system is clear.
Acknowledgements: Our interest in finding a characterization of $\Gamma$–shift-invariant closed subspaces originated with a question posed by Azita Mayeli. We also thank a referee of an earlier version of this paper who brought [3] to our attention.
2 Preliminaries
Let $n$ be a positive integer and let $\mathbb{R}^{n}$ denote Euclidean space with $x\in\mathbb{R}^{n}$ being considered as a column vector. For $1\leq p\leq\infty$, $L^{p}(\mathbb{R}^{n})$ denotes the standard Lebesgue space with respect to Lebesgue measure on $\mathbb{R}^{n}$ and $\|\cdot\|_{p}$ denotes the usual norm in $L^{p}(\mathbb{R}^{n})$. We use the following version of the Fourier transform. For $f\in L^{1}(\mathbb{R}^{n})$ and $\omega\in\mathbb{R}^{n}$, let $\widehat{f}(\omega)=\int_{\mathbb{R}^{n}}f(x)e^{2\pi i\omega\cdot x}dx$. By Plancherel’s Theorem, there is a unitary map $\mathcal{F}$ on $L^{2}(\mathbb{R}^{n})$ such that $\mathcal{F}f=\widehat{f}$, for all $f\in L^{1}(\mathbb{R}^{n})\cap L^{2}(\mathbb{R}^{n})$. We call both $\mathcal{F}$ and the map $f\mapsto\widehat{f}$ the Fourier transform. We will often use $\widehat{f}$ instead of $\mathcal{F}f$, for any $f\in L^{2}(\mathbb{R}^{n})$.
If we consider $\mathbb{R}^{n}$ as a topological group with addition of vectors as the group operation, a subgroup $\mathcal{L}$ of $\mathbb{R}^{n}$ is called a lattice in $\mathbb{R}^{n}$ if $\mathcal{L}$ is discrete and $\mathbb{R}^{n}/\mathcal{L}$ is compact. The dual lattice of $\mathcal{L}$ is $\mathcal{L}^{*}=\{\nu\in\mathbb{R}^{n}:\nu\cdot k\in\mathbb{Z}\}$. That is, $\nu\in\mathbb{R}^{n}$ is in $\mathcal{L}^{*}$ if and only if $e^{2\pi i\nu\cdot k}=1$, for all $k\in\mathcal{L}$. If $\mathcal{L}$ is a lattice in $\mathbb{R}^{n}$, then there exists an invertible $n\times n$ matrix $B$ such that $\mathcal{L}=B\mathbb{Z}^{n}=\{Bk:k\in\mathbb{Z}^{n}\}$. Then $\mathcal{L}^{*}=(B^{-1})^{t}\mathbb{Z}^{n}$, where $(B^{-1})^{t}$ is the transpose of the inverse of $B$. Let $Q=\{(B^{-1})^{t}\theta:\theta\in\left[-\frac{1}{2},\frac{1}{2}\right)^{n}\}$. Then $\mathbb{R}^{n}$ is the disjoint union of the $Q+\nu$, $\nu\in\mathcal{L}^{*}$. We equip $Q$ with the restriction of Lebesgue measure.
We will need a form of the Poisson summation formula. For $g\in L^{1}(\mathbb{R}^{n})$, the periodization $\mathcal{P}_{\mathcal{L}^{*}}g$ of $g$ with respect to $\mathcal{L}^{*}$ is given by
$\mathcal{P}_{\mathcal{L}^{*}}g(\theta)=\sum_{\nu\in\mathcal{L}^{*}}g(\theta+\nu),$
for any $\theta\in\mathbb{R}^{n}$ for which the series converges.
Lemma 2.1.
Let $\mathcal{L}$ be a lattice in $\mathbb{R}^{n}$ and let $\mathcal{L}^{*}$ be its dual lattice. Let $g\in L^{1}(\mathbb{R}^{n})$. Then $\sum_{\nu\in\mathcal{L}^{*}}g(\theta+\nu)$, converges for a.e. $\theta\in\mathbb{R}^{n}$ and $\mathcal{P}_{\mathcal{L}^{*}}g\big|_{Q}\in L^{1}(Q)$.
Proof.
We adapt the standard argument as in the proof of Theorem 8.31 of [11].
$\int_{Q}\sum_{\nu\in\mathcal{L}^{*}}|g(\theta+\nu)|d\theta$ $=$ $\sum_{\nu\in\mathcal{L}^{*}}\int_{Q}|g(\theta+\nu)|d\theta$
$=$ $\sum_{\nu\in\mathcal{L}^{*}}\int_{Q+\nu}|g(\theta)|d\theta=\int_{\mathbb{R}^{n}}|g(\theta)|d\theta=\|g\|_{1}.$
By the sum variation of the Dominated Convergence Theorem, $\sum_{\nu\in\mathcal{L}^{*}}g(\theta+\nu)$ converges for a.e. $\theta\in Q$ and $\mathcal{P}_{\mathcal{L}^{*}}g\big|_{Q}\in L^{1}(Q)$. Since $\mathcal{P}_{\mathcal{L}^{*}}g$ is $\mathcal{L}^{*}$-periodic, $\sum_{\nu\in\mathcal{L}^{*}}g(\theta+\nu)$ converges for a.e. $\theta\in\mathbb{R}^{n}$. ∎
Proposition 2.2.
Let $\mathcal{L}$ be a lattice in $\mathbb{R}^{n}$ and let $\mathcal{L}^{*}$ be its dual lattice. Let $g\in L^{1}(\mathbb{R}^{n})$. For each $\ell\in\mathcal{L}$,
[…]
Moreover, if $\widehat{g}(\ell)=0$, for all $\ell\in\mathcal{L}$, then $\mathcal{P}_{\mathcal{L}^{*}}g=0$.
Proof.
Again, dominated convergence justifies the interchange of sum and integral below. For $\ell\in\mathcal{L}$,
$\int_{Q}\mathcal{P}_{\mathcal{L}^{*}}g(\theta)e^{2\pi i\theta\cdot\ell}d\theta$ $=$ $\int_{Q}\sum_{\nu\in\mathcal{L}^{*}}g(\theta+\nu)e^{2\pi i\theta\cdot\ell}d\theta=\sum_{\nu\in\mathcal{L}^{*}}\int_{Q}g(\theta+\nu)e^{2\pi i\theta\cdot\ell}d\theta$
$=$ $\sum_{\nu\in\mathcal{L}^{*}}\int_{Q+\nu}g(\theta)e^{2\pi i(\theta-\nu)\cdot\ell}d\theta=\sum_{\nu\in\mathcal{L}^{*}}\int_{Q+\nu}g(\theta)e^{2\pi i\theta\cdot\ell}d\theta$
[…]
For the last statement in the proposition, we do a linear change of variables. Let $C=(B^{-1})^{t}$, where $\mathcal{L}=B\mathbb{Z}^{n}$. That is, $C$ is the invertible matrix such that $\mathcal{L}^{*}=C\mathbb{Z}^{n}$. Now define $h\in L^{1}(\mathbb{R}^{n})$ by $h(\omega)=g(C\omega)$, for a.e. $\omega\in\mathbb{R}^{n}$. Form the $\mathbb{Z}^{n}$ periodization of $h$. That is,
$\mathcal{P}_{\mathbb{Z}^{n}}h(\omega)=\sum_{k\in\mathbb{Z}^{n}}h(\omega+k)=\sum_{k\in\mathbb{Z}^{n}}g(C\omega+Ck)=\mathcal{P}_{\mathcal{L}^{*}}g(C\omega),$
which converges for a.e. $\omega\in\mathbb{R}^{n}$, and $\mathcal{P}_{\mathbb{Z}^{n}}h$ is integrable over $\left[-\frac{1}{2},\frac{1}{2}\right)^{n}$. Let $(c_{j})_{j\in\mathbb{Z}^{n}}$ be the Fourier multi-series of the periodic function $\mathcal{P}_{\mathbb{Z}^{n}}h$. For each $j\in\mathbb{Z}^{n}$, compute the Fourier coefficient $c_{j}$ using the change of variables $\theta=C\omega$ and noting that $C^{-1}=B^{t}$. So
$c_{j}$ $=$ $\int_{\left[-\frac{1}{2},\frac{1}{2}\right)^{n}}\mathcal{P}_{\mathbb{Z}^{n}}h(\omega)e^{2\pi i\omega\cdot j}d\omega=\int_{\left[-\frac{1}{2},\frac{1}{2}\right)^{n}}\mathcal{P}_{\mathcal{L}^{*}}g(C\omega)e^{2\pi i\omega\cdot j}d\omega$
$=$ $|\det(B)|\int_{Q}\mathcal{P}_{\mathcal{L}^{*}}g(\theta)e^{2\pi i\theta\cdot Bj}d\theta=|\det(B)|\widehat{g}(Bj).$
Now $Bj\in\mathcal{L}^{*}$, for all $j\in\mathbb{Z}^{n}$. So, if $\widehat{g}(\ell)=0$, for all $\ell\in\mathcal{L}^{*}$, then every Fourier coefficient of $\mathcal{P}_{\mathbb{Z}^{n}}h$ is 0. This implies $\mathcal{P}_{\mathbb{Z}^{n}}h=0$, which implies $\mathcal{P}_{\mathcal{L}^{*}}g=0$. ∎
Corollary 2.3.
With the notation of Proposition 2.2, let $Y$ be an open subset of $\mathbb{R}^{n}$ such that $(Y+\nu)\cap(Y+\kappa)=\emptyset$ if $\nu,\kappa\in\mathcal{L}^{*}$ with $\nu\neq\kappa$. Let $g\in L^{1}(\mathbb{R}^{n})$ be such that $g(\theta)=0$, for a.e. $\theta\in\mathbb{R}^{n}\setminus Y$. If $\widehat{g}(\ell)=0$, for all $\ell\in\mathcal{L}$, then $g=0$.
Proof.
Under the assumption on $g$, $\mathcal{P}_{\mathcal{L}^{*}}g$ agrees with $g$ on $Y$, so Proposition 2.2 implies $g=0$. ∎
Remark 2.4.
For each $\ell\in\mathcal{L}$, let $e_{\ell}(\theta)=e^{2\pi i\theta\cdot\ell}$, for all $\theta\in\mathbb{R}^{n}$. In the language of [7], Definition 2.2, $\mathcal{D}=\{e_{\ell}|_{{}_{Y}}:\ell\in\mathcal{L}\}$ is a determining set for $L^{1}(Y)$.
The concept of a range function goes back to [15] and was used in [6] for the characterization of closed subspaces of $L^{2}(\mathbb{R}^{n})$ that are shift-invariant. The treatment of range functions given in [7] is most useful for our purposes. For a Hilbert space $\mathcal{H}$, let $Gr(\mathcal{H})$ denote the set of all closed subspaces of $\mathcal{H}$, the Grassmannian of $\mathcal{H}$. If $\mathcal{K}\in Gr(\mathcal{H})$, let $P_{\mathcal{K}}$ denote the orthogonal projection of $\mathcal{H}$ onto $\mathcal{K}$.
Definition 2.5.
Let $\mathcal{H}$ be a separable Hilbert space and let $(\mathcal{X},\Sigma)$ be a measurable space. A measurable range function for $\mathcal{H}$ based on $\mathcal{X}$, is a mapping $J:\mathcal{X}\to Gr(\mathcal{H})$ such that, for any $\xi,\eta\in\mathcal{H}$, the map $x\to\langle P_{J(x)}\xi,\eta\rangle$ is measurable.
In all the cases we consider here, $\mathcal{H}$ is a separable Hilbert space. A map $F:\mathcal{X}\to\mathcal{H}$ is measurable if $x\to\langle F(x),\eta\rangle$ is a measurable complex-valued function, for each $\eta\in\mathcal{H}$. If $\mu$ is a positive measure on $(\mathcal{X},\Sigma)$, let
$L^{2}(\mathcal{X},\mu,\mathcal{H})=\left\{F:\mathcal{X}\to\mathcal{H}\big|F\text{ is measurable and }\int_{\mathcal{X}}\|F(x)\|^{2}d\mu(x)<\infty\right\},$
with the usual identification of functions that agree $\mu$-almost everywhere. If the measure $\mu$ is understood from context, we write $L^{2}(\mathcal{X},\mathcal{H})$. The inner product of $F_{1},F_{2}\in L^{2}(\mathcal{X},\mathcal{H})$ is given by
$\langle F_{1},F_{2}\rangle=\int_{\mathcal{X}}\langle F_{1}(x),F_{2}(x)\rangle d\mu(x).$
Let $J$ be a measurable range function for $\mathcal{H}$ based on $\mathcal{X}$. Then, for any $F\in L^{2}(\mathcal{X},\mathcal{H})$, $x\to\langle\mathcal{P}_{J(x)}F(x),\eta\rangle$ is measurable, for every $\eta\in\mathcal{H}$. Given $J$, define $\mathcal{M}_{J}=\{F\in L^{2}(\mathcal{X},\mathcal{H}):F(x)\in J(x),\text{ for }\mu\text{-a.e. }x\in\mathcal{X}\}$. The next Proposition gathers together Propositions 2.1, 2.2, and 2.3 of [7].
Proposition 2.6.
Let JJand $K$ be measurable range functions for $\mathcal{H}$ based on $\mathcal{X}$. Then
(i) $\mathcal{M}_{J}$ is a closed subspace of $L^{2}(\mathcal{X},\mathcal{H})$,
(ii) if $\mathcal{P}_{\mathcal{M}_{J}}$ is the orthogonal projection of $L^{2}(\mathcal{X},\mathcal{H})$ onto $\mathcal{M}_{J}$, then, for any $F\in L^{2}(\mathcal{X},\mathcal{H})$, $\left(\mathcal{P}_{\mathcal{M}_{J}}F\right)(x)=\mathcal{P}_{J(x)}F(x)$, for a.e. $x\in\mathcal{X}$, and
(iii) $\mathcal{M}_{J}=\mathcal{M}_{K}$ if and only if $J(x)=K(x)$, for a.e. $x\in\mathcal{X}$.
If $\mathcal{H}$ is a Hilbert space, let $\mathcal{U}(\mathcal{H})$ denote the group of all unitary operators on $\mathcal{H}$. If $\Lambda$ is any discrete group, a unitary representation of $\Lambda$ on $\mathcal{H}$ is a homomorphism $\sigma:\Lambda\to\mathcal{U}(\mathcal{H})$. Let $\sigma_{1}$ and $\sigma_{2}$ be unitary representations of $\Lambda$ on $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$, respectively. If there exists a unitary map $W:\mathcal{H}_{1}\to\mathcal{H}_{2}$ such that $W\sigma_{1}(a)=\sigma_{2}(a)W$, for all $a\in\Lambda$, then $\sigma_{1}$ and $\sigma_{2}$ are called equivalent.
3 Crystal groups
Let ${\rm O}_{n}$ denote the compact group of orthogonal $n\times n$ real matrices. For $x\in\mathbb{R}^{n}$ and $A\in{\rm O}_{n}$, define $[x,A]$ to be the affine map, $z\to A(z+x)$, of $\mathbb{R}^{n}$. Let
${\rm Iso}_{n}(\mathbb{R})=\{[x,A]:x\in\mathbb{R}^{n},A\in{\rm O}_{n}\},$
the group of all isometries of $\mathbb{R}^{n}$. The composition of isometries is the group product: For $x,A],[y,B]\in{\rm Iso}_{n}(\mathbb{R})$, $x,A][y,B]=[B^{-1}x+y,AB]$ and $x,A]^{-1}=[-Ax,A^{-1}]$. Note that $0,{\rm id}]$ is the identity element of ${\rm Iso}_{n}(\mathbb{R})$, where $\rm id$} denotes the identity $n\times n$ matrix. Given the product topology of $\mathbb{R}^{n}\times{\rm O}_{n}$, ${\rm Iso}_{n}(\mathbb{R})$ is a locally compact group. Let ${\rm Trans}_{n}=\{[x,{\rm id}]:x\in\mathbb{R}^{n}\}$, the set of pure translations in ${\rm Iso}_{n}(\mathbb{R})$. This is a closed normal subgroup of ${\rm Iso}_{n}(\mathbb{R})$. Let $q:{\rm Iso}_{n}(\mathbb{R})\to{\rm O}_{n}$ be the homomorphism given by $q[x,A]=A$, for all $x,A]\in{\rm Iso}_{n}(\mathbb{R})$.
A crystal group is a discrete subgroup $\Gamma$ of ${\rm Iso}_{n}(\mathbb{R})$ such that $\mathbb{R}^{n}/\Gamma$ is compact, where $\mathbb{R}^{n}/\Gamma$ is the set of all $\Gamma$-orbits, with the quotient topology. Section 7.5 of [20] presents the basic properties of crystal groups. The translation subgroup of $\Gamma$ is $T=\Gamma\cap{\rm Trans}_{n}$, which is a normal subgroup, and the point group is $\Pi=q(\Gamma)$, which is isomorphic to the quotient group $\Gamma/T$. Then $\Pi$ is a finite subgroup of ${\rm O}_{n}$, while $T$ is a free abelian group of rank $n$. Indeed, $\mathcal{L}=\{\ell\in\mathbb{R}^{n}:[\ell,{\rm id}]\in T\}$ is a lattice in $\mathbb{R}^{n}$. For $\ell,{\rm id}]\in T$ and $x,M]\in\Gamma$,
$x,M][\ell,{\rm id}][x,M]^{-1}=[x+\ell,M][-Mx,M^{-1}]=[M\ell,{\rm id}].$
This shows that if $\ell\in\mathcal{L}$, then $M\ell\in\mathcal{L}$, for all $M\in\Pi$.
Fix a cross-section $\gamma:\Pi\to\Gamma$ of the $T$-cosets in $\Gamma$; so $q\big(\gamma(M)\big)=M$, for all $M\in\Pi$. With $\gamma$ fixed, for each $M\in\Pi$, let $x_{M}\in\mathbb{R}^{n}$ be such that $\gamma(M)=[x_{M},M]$. Then $\Gamma=\{[\ell+x_{M},M]:\ell\in\mathcal{L},M\in\Pi\}$. For many crystal groups, $0,M]\in\Gamma$, for all $M\in\Pi$. Then, choose $x_{M}=0$, for each $M\in\Pi$. When this can be done, $\Gamma$ is isomorphic to the semidirect product $\mathcal{L}\rtimes\Pi$. As pointed out in the introduction, such crystal groups are called symmorphic.
Definition 3.1.
Let $\Gamma$ be a crystal group in dimension $n$ with point group $\Pi$. A vector system associated with $\Gamma$ is a map $M\mapsto x_{M}$ of $\Pi$ into $\mathbb{R}^{n}$ such that $x_{M},M]\in\Gamma$, for all $M\in\Pi$.
Lemma 3.2.
If $\{x_{M}:M\in\Pi\}$ is a vector system associated with $\Gamma$, then $N^{-1}x_{K}+x_{N}-x_{KN}\in\mathcal{L}$, for all $K,N\in\Pi$.
Proof.
For $K,N\in\Pi$, calculate the following element of $\Gamma$:
$\begin{array}[]{ll}[x_{K},K][x_{N},N][x_{KN},KN]^{-1}&=[N^{-1}x_{K}+x_{N},KN][-KNx_{KN},(KN)^{-1}]\\ &=[Kx_{K}+KNx_{N}-KNx_{KN},{\rm id}]\in T.\end{array}$
Thus, $Kx_{K}+KNx_{N}-KNx_{KN}\in\mathcal{L}$. Then
$N^{-1}x_{K}+x_{N}-x_{KN}=N^{-1}K^{-1}\big(Kx_{K}+KNx_{N}-KNx_{KN}\big)$
is in $\mathcal{L}$ as well. ∎
[…]
$(\kappa,L)(\nu,M)=(M^{-1}\kappa+\nu,LM),\text{ for }(\kappa,L),(\nu,M)\in\mathcal{L}^{*}\rtimes\Pi.$
This auxiliary group might be isomorphic to $\Gamma$, but this is not always the case. As with $\Gamma$, $\mathcal{L}^{*}\rtimes\Pi$ acts on $\mathbb{R}^{n}$. For $(\nu,M)\in\mathcal{L}^{*}\rtimes\Pi$ and $\omega\in\mathbb{R}^{n}$, let $(\nu,M)\cdot\omega=M(\omega+\nu)$. This identifies $\mathcal{L}^{*}\rtimes\Pi$ with a discrete group of isometries of $\mathbb{R}^{n}$ such that, since $\mathcal{L}^{*}$ is a full-rank lattice, $\mathbb{R}^{n}/(\mathcal{L}^{*}\rtimes\Pi)$ is compact. That is, $\mathcal{L}^{*}\rtimes\Pi$ is also a crystal group. Let $\Gamma^{*}=\mathcal{L}^{*}\rtimes\Pi$.
For any $\omega\in\mathbb{R}^{n}$, let $\Gamma^{*}\omega=\{M(\omega+\nu):(\nu,M)\in\Gamma^{*}\}$, the $\Gamma^{*}$-orbit of $\omega$, and let $\Gamma^{*}_{\omega}=\{(\nu,M)\in\Gamma^{*}:M(\omega+\nu)=\omega\}$, the stabilizer of $\omega$. There exist points $\omega$ in $\mathbb{R}^{n}$ such that $\Gamma^{*}_{\omega}=\{(0,{\rm id})\}$ (see Theorem 6.6.12 of [20]). Fix $\omega_{0}\in\mathbb{R}^{n}$ such that $\Gamma^{*}_{\omega_{0}}=\{(0,{\rm id})\}$. For each $(\nu,M)\in\Gamma^{*}\setminus\{(0,{\rm id})\}$, let $H_{(\nu,M)}=\{\omega\in\mathbb{R}^{n}:\|\omega-\omega_{0}\|<\|\omega-M(\omega_{0}+\nu)\|\}$.
Definition 3.3.
The Dirichlet domain for $\Gamma^{*}$ containing $\omega_{0}$ is
$\Omega_{\omega_{0}}=\bigcap\{H_{(\nu,M)}:(\nu,M)\in\Gamma^{*},(\nu,M)\neq(0,{\rm id})\}.$
Let $A\subseteq\mathbb{R}^{n}$, then $A+\nu=\{\omega+\nu:\omega\in A\}$, for $\nu\in\mathcal{L}^{*}$, and $MA=\{M\omega:\omega\in A\}$, for $M\in\Pi$. Also, $\partial A$ denotes the boundary of $A$.
Proposition 3.4.
Let $\omega_{0}\in\mathbb{R}^{n}$ be such that $\Gamma^{*}_{\omega_{0}}=\{(0,{\rm id})\}$ and let $\Omega=\Omega_{\omega_{0}}$. Then $\Omega$ has the following properties:
  1. 1.
    $\Omega$ is open
  2. 2.
    $\Omega$ is convex
  3. 3.
    For $(\kappa,L),(\nu,M)\in\Gamma^{*},(\kappa,L)\neq(\nu,M)$, $(L(\Omega+\kappa))\cap(M(\Omega+\nu))=\emptyset$
  4. 4.
    $\cup_{(\nu,M)\in\Gamma^{*}}M(\overline{\Omega}+\nu)=\mathbb{R}^{n}$
  5. 5.
    $\cup_{(\nu,M)\in\Gamma^{*}}M(\Omega+\nu)$ is a co-null subset of $\mathbb{R}^{n}$.
Proof.
Properties 1, 2, 3, and 4 are well-known (see, for example, Theorem 6.6.13 and the definition of fundamental domain on page 233 of [20]). Since $\Omega$ is open and convex $\partial\Omega$ has Lebesgue measure 0. Each map $\omega\to M(\omega+\nu)$ is an isometry, so $\partial(M(\Omega+\nu))=M(\partial\Omega+\nu)$ is a null set, for each $(\nu,M)\in\Gamma^{*}$. Now $\mathbb{R}^{n}\setminus\cup_{(\nu,M)\in\Gamma^{*}}M(\Omega+\nu)\subseteq\cup_{(\nu,M)\in\Gamma^{*}}M(\partial\Omega+\nu)$, which is a null set since $\Gamma^{*}$ is countable. This implies 5. ∎
Let $\Pi\Omega=\cup_{M\in\Pi}M\Omega$. Then $\Pi\Omega$ is a fundamental domain for $\mathcal{L}^{*}$. In particular, we have the following corollary.
Corollary 3.5.
If $\nu_{1},\nu_{2}\in\mathcal{L}^{*}$, $\nu_{1}\neq\nu_{2}$, then $(\Pi\Omega+\nu_{1})\cap(\Pi\Omega+\nu_{2})=\emptyset$ and $\cup_{\nu\in\mathcal{L}^{*}}(\Pi\Omega+\nu)$ is co-null in $\mathbb{R}^{n}$.
4 Several Unitary Transformations
Throughout this section, $\Gamma$ is a fixed crystal group with $\mathcal{L}$, $\mathcal{L}^{*}$, $\Pi$, $T$, $\Omega$, and a vector system $\{x_{M}:M\in\Pi\}$ as in Section 3. In order to illuminate aspects of the natural unitary representation of the crystal group $\Gamma$ in the next section, the Hilbert space $L^{2}(\mathbb{R}^{n})$ will be transformed into a number of different realizations by specific unitary maps. The first is the Fourier transform viewed as a unitary map of $L^{2}(\mathbb{R}^{n})$ onto itself.
For the other Hilbert spaces that arise, we equip the open subsets $\Omega$ and $\Pi\Omega$ with the restriction of Lebesgue measure and $\Omega\times\Pi\times\mathcal{L}^{*}$ with the product measure, where $\Pi$ and $\mathcal{L}^{*}$ have counting measure. We can use $\Omega\times\Pi\times\mathcal{L}^{*}$ to almost parametrize $\mathbb{R}^{n}$. Let $X=\cup_{(L,\nu)\in\Pi\times\mathcal{L}^{*}}L(\Omega+\nu)$. By Proposition 3.4, $X$ is an open co-null subset of $\mathbb{R}^{n}$.
Lemma 4.1.
The map $\phi:\Omega\times\Pi\times\mathcal{L}^{*}\to X$ given by $\phi(\omega,L,\nu)=L(\omega+\nu)$ is a measure preserving homeomorphism.
Proof.
Routine. ∎
For any $\theta\in X$, let $\omega_{\theta}\in\Omega$, $L_{\theta}\in\Pi$, and $\nu_{\theta}\in\mathcal{L}^{*}$ be such that $L_{\theta}(\omega_{\theta}+\nu_{\theta})=\theta$. For each $\xi\in L^{2}(\mathbb{R}^{n})$, let $W_{1}\xi=\xi\circ\phi$. Then $W_{1}\xi$ is a measurable function on $\Omega\times\Pi\times\mathcal{L}^{*}$.
Lemma 4.2.
The map $W_{1}$ is a unitary from $L^{2}(\mathbb{R}^{n})$ onto $L^{2}(\Omega\times\Pi\times\mathcal{L}^{*})$ and $W_{1}^{-1}$ is given by, for $g\in L^{2}(\Omega\times\Pi\times\mathcal{L}^{*})$, $W_{1}^{-1}g(\theta)=g(\omega_{\theta},L_{\theta},\nu_{\theta})$, for a.e. $\theta\in X.$
Proof.
Also routine. ∎
Our next unitary transformation is a simple unitary operator on the Hilbert space $L^{2}(\Omega\times\Pi\times\mathcal{L}^{*})$. Recall that, for each $L\in\Pi$, $x_{L}\in\mathbb{R}^{n}$ was selected so that $x_{L},L]\in\Gamma$. Define a function $w_{2}:\Omega\times\Pi\times\mathcal{L}^{*}\to\mathbb{T}$ by $w_{2}(\omega,L,\nu)=e^{-2\pi i\nu\cdot x_{L}}$, for all $(\omega,L,\nu)\in\Omega\times\Pi\times\mathcal{L}^{*}$. Pointwise multiplication by such a $\mathbb{T}$-valued continuous function is a unitary operator and we formulate this in a lemma.
Lemma 4.3.
For each $g\in L^{2}(\Omega\times\Pi\times\mathcal{L}^{*})$, let $W_{2}g=w_{2}g$. That is,
$(W_{2}g)(\omega,L,\nu)=e^{-2\pi i\nu\cdot x_{L}}g(\omega,L,\nu),$
for a.e. $(\omega,L,\nu)\in\Omega\times\Pi\times\mathcal{L}^{*}$. Then $W_{2}$ is a unitary operator on $L^{2}(\Omega\times\Pi\times\mathcal{L}^{*})$ and $W_{2}^{-1}g=\overline{w_{2}}g$, for all $g\in L^{2}(\Omega\times\Pi\times\mathcal{L}^{*})$.
Remark 4.4.
The function $w_{2}$ plays an important role in the characterization of $\Gamma$–shift-invariant closed subspaces of $L^{2}(\mathbb{R}^{n})$ obtained in Section 6, so it is important to note that this function is independent of the choice of the vector system associated with $\Gamma$. Indeed, if $\{x^{\prime}_{L}:L\in\Pi\}$ is another vector system associated with $\Gamma$, then, for each $L\in\Pi$, $x_{L},L]^{-1}[x^{\prime}_{L},L]=[x^{\prime}_{L}-x_{L},{\rm id}]\in T$, so $k=x^{\prime}_{L}-x_{L}\in\mathcal{L}$. Thus, $e^{-2\pi i\nu\cdot x^{\prime}_{L}}=e^{-2\pi i\nu\cdot(x_{L}+k)}=e^{-2\pi i\nu\cdot x_{L}}$, for all $\nu\in\mathcal{L}^{*}$.
Let $\{\delta_{\nu}:\nu\in\mathcal{L}^{*}\}$ be the orthonormal basis of $\ell^{2}(\mathcal{L}^{*})$ consisting of the usual delta functions. We will use $L^{2}\big(\Pi\Omega,\ell^{2}(\mathcal{L}^{*})\big)$, which is a separable Hilbert space. As we work with each of $L^{2}(\mathbb{R}^{n})$, $L^{2}(\Omega\times\Pi\times\mathcal{L}^{*})$, $\ell^{2}(\mathcal{L}^{*})$, and $L^{2}\big(\Pi\Omega,\ell^{2}(\mathcal{L}^{*})\big)$, any inner product that arises will be denoted $\langle\cdot,\cdot\rangle$ and the corresponding norm denoted by $\|\cdot\|$, relying on the reader to know which Hilbert space this inner product or norm belongs to from the context.
We now turn to the definition of a unitary transformation of $L^{2}(\Omega\times\Pi\times\mathcal{L}^{*})$ into $L^{2}\big(\Pi\Omega,\ell^{2}(\mathcal{L}^{*})\big)$. As above, for any $\theta\in\Pi\Omega$, there are unique $\omega_{\theta}\in\Omega$ and $L_{\theta}\in\Pi$ so that $L_{\theta}\omega_{\theta}=\theta$. For any $g\in L^{2}(\Omega\times\Pi\times\mathcal{L}^{*})$ and $\nu\in\mathcal{L}^{*}$, let $g_{\nu}(\theta)=g(\omega_{\theta},L_{\theta},\nu)$, for a.e. $\theta\in\Pi\Omega$. Then $g_{\nu}\in L^{2}(\Pi\Omega)$, for each $\nu\in\mathcal{L}^{*}$, and $\|g\|^{2}=\sum_{\nu\in\mathcal{L}^{*}}\|g_{\nu}\|^{2}$. Define $W_{3}g:\Pi\Omega\to\ell^{2}(\mathcal{L}^{*})$ by
$W_{3}g(\theta)=\sum_{\nu\in\mathcal{L}^{*}}g_{\nu}(\theta)\delta_{\nu},\text{ for a.e. }\theta\in\Pi\Omega.$ (1)
Lemma 4.5.
If $W_{3}g$ is defined by (1), then $W_{3}g\in L^{2}\big(\Pi\Omega,\ell^{2}(\mathcal{L}^{*})\big)$, for each $g\in L^{2}(\Omega\times\Pi\times\mathcal{L}^{*})$ and $W_{3}$ is a unitary map of $L^{2}(\Omega\times\Pi\times\mathcal{L}^{*})$ onto $L^{2}\big(\Pi\Omega,\ell^{2}(\mathcal{L}^{*})\big)$. Moreover, for $F\in L^{2}\big(\Pi\Omega,\ell^{2}(\mathcal{L}^{*})\big)$,
$W_{3}^{-1}F(\omega,L,\nu)=\langle F(L\omega),\delta_{\nu}\rangle,$ (2)
for a.e. $(\omega,L,\nu)\in\Omega\times\Pi\times\mathcal{L}^{*}$.
Proof.
For each $\kappa\in\mathcal{L}^{*}$, the map $\theta\mapsto\langle W_{3}g(\theta),\delta_{\kappa}\rangle=g_{\kappa}(\theta)$ is measurable. Since $\{\delta_{\kappa}:\kappa\in\mathcal{L}^{*}\}$ is a countable basis of $\ell^{2}(\mathcal{L}^{*})$, $W_{3}g$ is measurable. Moreover, using (1)
$\int_{\Pi\Omega}\|W_{3}g(\theta)\|^{2}d\theta=\int_{\Pi\Omega}\sum_{\nu\in\mathcal{L}^{*}}|g_{\nu}(\theta)|^{2}d\theta=\sum_{\nu\in\mathcal{L}^{*}}\int_{\Pi\Omega}|g_{\nu}(\theta)|^{2}d\theta=\|g\|^{2}.$ (3)
Therefore, $W_{3}g\in L^{2}\big(\Pi\Omega,\ell^{2}(\mathcal{L}^{*})\big)$ and $W_{3}$ is clearly linear. By (3), $W_{3}$ is an isometry. If $F\in L^{2}\big(\Pi\Omega,\ell^{2}(\mathcal{L}^{*})\big)$ and (2) is used to define $W_{3}^{-1}F$ on $\Omega\times\Pi\times\mathcal{L}^{*}$, then one verifies, with an argument similar to (3) in reverse, that $W_{3}^{-1}F$ is square-integrable on $\Omega\times\Pi\times\mathcal{L}^{*}$ and $W_{3}W_{3}^{-1}F=F$. ∎
5 The natural representation
We continue with the notation of the previous section. The action of $\Gamma$ on $\mathbb{R}^{n}$ generates a unitary representation. For $x,M]\in\Gamma$, $\pi[x,M]$ is the unitary operator given by, for $f\in L^{2}(\mathbb{R}^{n})$,
$\pi[x,M]f(z)=f([x,M]^{-1}z)=f(M^{-1}z-x),\text{ for a.e. }z\in\mathbb{R}^{n}.$
It is routine to show that $\pi$ is a homomorphism of $\Gamma$ into $\mathcal{U}\big(L^{2}(\mathbb{R}^{n})\big)$. As mentioned in the introduction, we will call $\pi$ the natural representation of $\Gamma$. The Fourier transform, as a unitary map $\mathcal{F}$ on $L^{2}(\mathbb{R}^{n})$, intertwines $\pi$ with an equivalent representation $\widehat{\pi}$ by $\widehat{\pi}[\ell,M]=\mathcal{F}\pi[\ell,M]\mathcal{F}^{-1}$. A brief calculation, using the fact that $M^{-1}$ agrees with the transpose matrix of $M\in{\rm O}_{n}$, shows that, for $\xi\in L^{2}(\mathbb{R}^{n})$,
$\widehat{\pi}[x,M]\xi(\omega)=e^{2\pi i(M^{-1}\omega)\cdot x}\xi(M^{-1}\omega),\text{ for a.e. }\omega\in\mathbb{R}^{n}.$ (4)
Recall from Lemma 4.2 that $W_{1}\xi(\omega,L,\nu)=\xi\big(L(\omega+\nu)\big)$, for almost every $(\omega,L,\nu)\in\Omega\times\Pi\times\mathcal{L}^{*}$, for each $\xi\in L^{2}(\mathbb{R}^{n})$, defines a unitary map of $L^{2}(\mathbb{R}^{n})$ onto $L^{2}(\Omega\times\Pi\times\mathcal{L}^{*})$. Let $\widehat{\pi}_{1}$ be the representation of $\Gamma$ on $L^{2}(\Omega\times\Pi\times\mathcal{L}^{*})$ given by $\widehat{\pi}_{1}[x,M]=W_{1}\widehat{\pi}[x,M]W_{1}^{-1}$, for $x,M]\in\Gamma$. For $g\in L^{2}(\Omega\times\Pi\times\mathcal{L}^{*})$, let $\xi=W_{1}^{-1}g$. Then, for $x,M]\in\Gamma$ and a.e. $\omega\in\Omega$, $L\in\Pi$, and $\nu\in\mathcal{L}^{*}$.
$\widehat{\pi}_{1}[x,M]g(\omega,L,\nu)=\widehat{\pi}[x,M]\xi\big(L(\omega+\nu)\big)=e^{2\pi i(M^{-1}L(\omega+\nu))\cdot x}\xi\big(M^{-1}L(\omega+\nu)\big).$
Setting $\theta=M^{-1}L(\omega+\nu)$, note that $\omega_{\theta}=\omega$, $L_{\theta}=M^{-1}L$, and $\nu_{\theta}=\nu$. Thus,
$\widehat{\pi}_{1}[x,M]g(\omega,L,\nu)=e^{2\pi i(M^{-1}L(\omega+\nu))\cdot x}g(\omega,M^{-1}L,\nu),$
for a.e. $(\omega,L,\nu)\in\Omega\times\Pi\times\mathcal{L}^{*}$. Since $x,M]\in\Gamma$, there exists $\ell\in\mathcal{L}$ so that $x=\ell+x_{M}$. For any $\nu\in\mathcal{L}^{*}$ and $N\in\Pi$, $e^{2\pi i(N\nu)\cdot\ell}=1$. Therefore, $e^{2\pi i(M^{-1}L(\omega+\nu))\cdot x}$ can be written as $e^{2\pi i(M^{-1}L\nu)\cdot x_{M}}e^{2\pi i(M^{-1}L\omega)\cdot x}=e^{2\pi i\nu\cdot(L^{-1}Mx_{M})}e^{2\pi i(M^{-1}L\omega)\cdot x}$ and
$\widehat{\pi}_{1}[x,M]g(\omega,L,\nu)=e^{2\pi i\nu\cdot(L^{-1}Mx_{M})}e^{2\pi i(M^{-1}L\omega)\cdot x}g(\omega,M^{-1}L,\nu).$ (5)
We can use the rather elementary, but strategic, unitary $W_{2}$ from Lemma 4.3 to eliminate the first factor in the right hand side of (5). Recall that $W_{2}g=w_{2}g$, for all $g\in L^{2}(\Omega\times\Pi\times\mathcal{L}^{*})$, where $w_{2}(\omega,L,\nu)=e^{-2\pi i\nu\cdot x_{L}}$, for $(\omega,L,\nu)\in\Omega\times\Pi\times\mathcal{L}^{*}$. Let $\widehat{\pi}_{2}[x,M]=W_{2}\widehat{\pi}_{1}[x,M]W_{2}^{-1}$, for all $x,M]\in\Gamma$. Then, for $g\in L^{2}(\Omega\times\Pi\times\mathcal{L}^{*})$,
$\widehat{\pi}_{2}[x,M]g(\omega,L,\nu)=e^{-2\pi i\nu\cdot x_{L}}\widehat{\pi}_{1}[x,M]W_{2}^{-1}g(\omega,L,\nu)$
$=e^{-2\pi i\nu\cdot x_{L}}e^{2\pi i\nu\cdot(L^{-1}Mx_{M})}e^{2\pi i(M^{-1}L\omega)\cdot x}W_{2}^{-1}g(\omega,M^{-1}L,\nu)$
$=e^{-2\pi i\nu\cdot x_{L}}e^{2\pi i\nu\cdot(L^{-1}Mx_{M})}e^{2\pi i(M^{-1}L\omega)\cdot x}e^{2\pi i\nu\cdot x_{M^{-1}L}}g(\omega,M^{-1}L,\nu).$
The scalar in the previous line can be rearranged to
$e^{2\pi i\nu\cdot(L^{-1}Mx_{M}+x_{M^{-1}L}-x_{L})}e^{2\pi i(L\omega)\cdot(Mx)}.$
Consider the expression $L^{-1}Mx_{M}+x_{M^{-1}L}-x_{L}$. By Lemma 3.2 with $K=M$ and $N=M^{-1}L$, we have that $L^{-1}Mx_{M}+x_{M^{-1}L}-x_{L}=\ell$, for some $\ell\in\mathcal{L}$. Since $\nu\in\mathcal{L}^{*}$, $e^{2\pi i\nu\cdot(L^{-1}Mx_{M}+x_{M^{-1}L}-x_{L})}=1$. Finally, for $x,M]\in\Gamma$ and $g\in L^{2}(\Omega\times\Pi\times\mathcal{L}^{*})$,
$\widehat{\pi}_{2}[x,M]g(\omega,L,\nu)=e^{2\pi i(L\omega)\cdot(Mx)}g(\omega,M^{-1}L,\nu)$ (6)
for a.e. $\omega\in\Omega$, $L\in\Pi$ and $\nu\in\mathcal{L}^{*}$.
The next step is to conjugate by the unitary $W_{3}:L^{2}(\Omega\times\Pi\times\mathcal{L}^{*})\to L^{2}\big(\Pi\Omega,\ell^{2}(\mathcal{L}^{*})\big)$ from Lemma 4.5. Let $\widehat{\pi}_{3}[x,M]=W_{3}\widehat{\pi}_{2}[x,M]W_{3}^{-1}$, for all $x,M]\in\Gamma$. For $F\in L^{2}\big(\Pi\Omega,\ell^{2}(\mathcal{L}^{*})\big)$,
$\widehat{\pi}_{3}[x,M]F(\theta)$ $=$ $W_{3}\widehat{\pi}_{2}[x,M]W_{3}^{-1}F(\theta)$ (7)
$=$ $\sum_{\nu\in\mathcal{L}^{*}}\left(\widehat{\pi}_{2}[x,M]W_{3}^{-1}F(\omega_{\theta},L_{\theta},\nu)\right)\delta_{\nu}$
$=$ $e^{2\pi i(L_{\theta}\omega_{\theta})\cdot(Mx)}\sum_{\nu\in\mathcal{L}^{*}}W_{3}^{-1}F(\omega_{\theta},M^{-1}L_{\theta},\nu)\delta_{\nu}.$
Using (2) we get $\sum_{\nu\in\mathcal{L}^{*}}W_{3}^{-1}F(\omega_{\theta},M^{-1}L_{\theta},\nu)\delta_{\nu}=F(M^{-1}L_{\theta}\omega_{\theta})$. Recall that $L_{\theta}\omega_{\theta}=\theta$ and $M^{t}=M^{-1}$, so (7) implies
$\widehat{\pi}_{3}[x,M]F(\theta)=e^{2\pi i(M^{-1}\theta)\cdot x}F(M^{-1}\theta),\text{ for a.e. }\theta\in\Pi\Omega,$ (8)
and each $x,M]\in\Gamma$ and any $F\in L^{2}\big(\Pi\Omega,\ell^{2}(\mathcal{L}^{*})\big)$. We summarize these considerations in a proposition.
Proposition 5.1.
If $\widehat{\pi}_{2}[x,M]$ is defined by (6) and $\widehat{\pi}_{3}[x,M]$ is defined by (8), for each $x,M]\in\Gamma$, then $\widehat{\pi}_{2}$ and $\widehat{\pi}_{3}$ are unitary representations of $\Gamma$ that are each unitarily equivalent to the natural representation $\pi$.
6 Invariant closed subspaces
Definition 6.1.
Let $J:\Pi\Omega\to Gr\big(\ell^{2}(\mathcal{L}^{*})\big)$ be a measurable range function. We say $J$ is $\Pi$-invariant, if $J(L\theta)=J(\theta)$, for a.e. $\theta\in\Pi\Omega$ and every $L\in\Pi$.
Proposition 6.2.
Let $\mathcal{K}$ be a closed subspace of $L^{2}\big(\Pi\Omega,\ell^{2}(\mathcal{L}^{*})\big)$. Then $\mathcal{K}$ is $\widehat{\pi}_{3}$-invariant if and only if there exists a $\Pi$-invariant measurable range function, $J$, for $\ell^{2}(\mathcal{L}^{*})$ based on $\Pi\Omega$ such that $\mathcal{K}=\mathcal{M}_{J}$.
Proof.
First, suppose $J$ is a $\Pi$-invariant measurable range function for $\ell^{2}(\mathcal{L}^{*})$ based on $\Pi\Omega$ and $\mathcal{K}=\mathcal{M}_{J}$. Let $x,M]\in\Gamma$. For any $F\in\mathcal{M}_{J}$, we know that $F(\theta)\in J(\theta)$, for a.e. $\theta\in\Pi\Omega$, and $J$ is $\Pi$-invariant. This implies
$\widehat{\pi}_{3}[x,M]F(\theta)=e^{2\pi i\theta\cdot(Mx)}F(M^{-1}\theta)\in J(M^{-1}\theta)=J(\theta),$
for a.e. $\theta\in\Pi\Omega$. Thus, $\widehat{\pi}_{3}[x,M]F\in\mathcal{K}$, for all $F\in\mathcal{K}$ and $x,M]\in\Gamma$. That is, $\mathcal{K}$ is $\widehat{\pi}_{3}$-invariant.
Now suppose $\mathcal{K}$ is a $\widehat{\pi}_{3}$-invariant closed subspace of $L^{2}\big(\Pi\Omega,\ell^{2}(\mathcal{L}^{*})\big)$. By Remark 2.4, $\mathcal{D}=\{e_{\ell}|_{{}_{\Pi\Omega}}:\ell\in\mathcal{L}\}$ is a determining set for $L^{1}(\Pi\Omega)$, where $e_{\ell}(\theta)=e^{2\pi i\theta\cdot\ell}$, for all $\theta\in\mathbb{R}^{n}$. If $F\in\mathcal{K}$ and $\ell\in\mathcal{L}$, then $\ell,{\rm id}]\in T\subseteq\Gamma$ and (8) says
$\widehat{\pi}_{3}[\ell,{\rm id}]F(\theta)=e^{2\pi i\theta\cdot\ell}F(\theta)=e_{\ell}(\theta)F(\theta),\text{ for a.e. }\theta\in\Pi\Omega.$
Therefore, $e_{\ell}F\in\mathcal{K}$, for each $\ell\in\mathcal{L}$. In the language of [7], $\mathcal{K}$ is multiplicatively-invariant with respect to $\mathcal{D}$, denoted $\mathcal{D}-MI$. By Theorem 2.4 of [7], there exists a measurable range function $J:\Pi\Omega\to Gr\big(\ell^{2}(\mathcal{L}^{*})\big)$ so that
$\mathcal{K}=\mathcal{M}_{J}=\{F\in L^{2}\big(\Pi\Omega,\ell^{2}(\mathcal{L}^{*})\big):F(\theta)\in J(\theta),\text{ for a.e. }\theta\in\Pi\Omega\}.$
Moreover, $J$ is unique up to almost everywhere agreement and, if $\mathcal{A}$ is any countable dense subset of $\mathcal{K}$, $J(\theta)$ is the closed linear span of $\{\varphi(\theta):\varphi\in\mathcal{A}\}$, for a.e. $\theta\in\Pi\Omega$. Fix such an $\mathcal{A}$.
Fix $L\in\Pi$ so that $x_{L},L]\in\Gamma$ and use (8) again to get
$\widehat{\pi}_{3}[x_{L},L]F(L\omega)=e^{2\pi i\omega\cdot(x_{L})}F(\omega),\text{ for a.e. }\omega\in\Omega.$ (9)
For $\varphi\in\mathcal{A}$, let $\varphi^{L}=\widehat{\pi}_{3}[x_{L},L]\varphi$ and let $L\cdot\mathcal{A}=\{\varphi^{L}:\varphi\in\mathcal{A}\}$. Then $L\cdot\mathcal{A}$ is a dense subset of $\mathcal{K}$. Therefore, by (9), for a.e. $\omega\in\Omega$,
$J(L\omega)=\overline{\rm span}\{\varphi^{L}(L\omega):\varphi\in\mathcal{A}\}=\overline{\rm span}\{\varphi(\omega):\varphi\in\mathcal{A}\}=J(\omega).$
Since $L\in\Pi$ is arbitrary, $J$ is $\Pi$-invariant. ∎
Before formulating the main result, we recall the various ingredients: $\Gamma$ is a crystal group in dimension $n$ with point group $\Pi$ and associated lattice $\mathcal{L}$. The dual lattice of $\mathcal{L}$ is $\mathcal{L}^{*}$. There is an open subset $\Omega$ of $\mathbb{R}^{n}$ described in Proposition 3.4 and the collection of sets $\{L(\Omega+\nu):L\in\Pi,\nu\in\mathcal{L}^{*}\}$ tiles $\mathbb{R}^{n}$ up to a null set.
Theorem 6.3.
Let $\mathcal{V}$ be a closed subspace of $L^{2}(\mathbb{R}^{n})$. Then $\mathcal{V}$ is invariant under shifts by elements of $\Gamma$ if and only if there exists a $\Pi$-invariant measurable range function $J$ for $\ell^{2}(\mathcal{L}^{*})$ based on $\Pi\Omega$ such that, for any $f\in\mathcal{V}$, there exists $F\in\mathcal{M}_{J}$ so that, for almost every $\omega\in\Omega$, every $L\in\Pi$, and every $\nu\in\mathcal{L}^{*}$,
$\widehat{f}\big(L(\omega+\nu)\big)=e^{2\pi i\nu\cdot x_{L}}\langle F(L\omega),\delta_{\nu}\rangle.$ (10)
Proof.
Suppose that $\mathcal{V}$ is a closed subspace of $L^{2}(\mathbb{R}^{n})$ that is invariant under shifts by elements of $\Gamma$. That is, $\mathcal{V}$ is $\pi$-invariant. Let $U=W_{3}\circ W_{2}\circ W_{1}\circ\mathcal{F}$, a unitary map that intertwines the natural representation $\pi$ with $\widehat{\pi}_{3}$. Therefore, $U\mathcal{V}$ is a $\widehat{\pi}_{3}$-invariant closed subspace of $L^{2}\big(\Pi\Omega,\ell^{2}(\mathcal{L}^{*})\big)$. By Proposition 6.2, there exists a $\Pi$-invariant measurable range function $J:\Pi\Omega\to Gr\big(\ell^{2}(\mathcal{L}^{*})\big)$ such that $U\mathcal{V}=\mathcal{M}_{J}$. For any $f\in\mathcal{V}$, let $F=Uf$. This can be written as $W_{1}\widehat{f}=W_{2}^{-1}W_{3}^{-1}F$, which is an element of $L^{2}(\Omega\times\Pi\times\mathcal{L}^{*})$. By the definition of $W_{1}$ given just before Lemma 4.2,
$W_{1}\widehat{f}(\omega,L,\nu)=\widehat{f}\big(L(\omega+\nu)\big),$ (11)
for a.e. $(\omega,L,\nu)\in\Omega\times\Pi\times\mathcal{L}^{*}$. On the other hand, using the equations for $W_{2}^{-1}$ and $W_{3}^{-1}$ given in Lemmas 4.3 and 4.5, we have
$W_{2}^{-1}W_{3}^{-1}F(\omega,L,\nu)=e^{2\pi i\nu\cdot x_{L}}\langle F(L\omega),\delta_{\nu}\rangle,$ (12)
for a.e. $(\omega,L,\nu)\in\Omega\times\Pi\times\mathcal{L}^{*}$. Comparing (11) with (12) yields (10).
Conversely, suppose $J$ is a $\Pi$-invariant measurable range function for $\ell^{2}(\mathcal{L}^{*})$ based on $\Pi\Omega$ and that $\mathcal{V}$ is related to $\mathcal{M}_{J}$ via (10). Then $\mathcal{M}_{J}$ is $\widehat{\pi}_{3}$-invariant by Proposition 6.2 and unpacking (10) shows that $U^{-1}\mathcal{M}_{J}=\mathcal{V}$. Since $U$ is the intertwining unitary between $\pi$ and $\widehat{\pi}_{3}$, $\mathcal{V}$ must be $\pi$-invariant. Thus, $\mathcal{V}$ is invariant under shifts by elements of $\Gamma$. ∎
Theorem 6.3 is a direct generalization of Proposition 1.5 in [6], where it is attributed to Helson [15]. It is not surprising that the range function associated to a $\pi$-invariant closed subspace must be $\Pi$-invariant. However, the appearance of the factor $e^{2\pi i\nu\cdot x_{L}}$ in (10) is not so obvious. As noted earlier, for a given $L\in\Pi$, the choice of $x_{L}$ such that $x_{L},L]\in\Gamma$ is not unique. However, if $y\in\mathbb{R}^{n}$ also satisfies $y,L]\in\Gamma$, then $x_{L}-y\in\mathcal{L}$. Thus, $e^{2\pi i\nu\cdot y}=e^{2\pi i\nu\cdot x_{L}}$, for all $\nu\in\mathcal{L}^{*}$. To help understand the role this term plays, we present an example of a group, necessarily non-symmorphic, where this factor is nontrivial.
Example 6.4.
The patch of brick wall illustrated in Figure 1 is meant to be a region of a pattern extending in all directions. The symmetry group of this pattern is often denoted $pg$, so set $\Gamma=pg$. If the origin is placed at the bottom of one of the vertical line segments, let $\bf u$} denote a horizontal vector pointing right whose length equals the length of one brick. Let $\bf v$} be an upward pointing vector whose length equals twice the width of a brick. Then, we can take the translation lattice of $\Gamma$ to be $\mathcal{L}=\{j{\bf u}+k{\bf v}:j,k\in\mathbb{Z}\}$. Let $\sigma\in{\rm O}_{2}$ denote the reflection about the horizontal axis. This is not a symmetry of the pattern, but $\frac{1}{2}{\bf u},\sigma]\in\Gamma$. Indeed, the point group of $\Gamma$ is $\Pi=\{{\rm id},\sigma\}$, isomorphic to $\mathbb{Z}_{2}$. If $T=\{[j{\bf u}+k{\bf v},{\rm id}]:j,k\in\mathbb{Z}\}$ is the translation subgroup of $\Gamma$, then $\{{\rm id}\mapsto 0;\ \sigma\mapsto\frac{1}{2}{\bf u}\}$ is a vector system associated with $\Gamma$.
[…]
Figure 1: A pattern illustrating the symmetries of the wallpaper group $pg$.
Fix the basis for $\mathbb{R}^{2}$ to be $\{{\bf u},{\bf v}\}$ and represent elements $x\in\mathbb{R}^{2}$ as $x=\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}$. Then $\mathcal{L}=\left\{\ell=\begin{pmatrix}\ell_{1}\\ \ell_{2}\end{pmatrix}:\ell_{1},\ell_{2}\in\mathbb{Z}\right\}=\mathbb{Z}^{2}$, viewed as a lattice in $\mathbb{R}^{2}$. So $\mathcal{L}^{*}=\mathbb{Z}^{2}$, as well. We set $\Gamma^{*}=\mathcal{L}^{*}\rtimes\Pi$. Let $\omega_{0}=\begin{pmatrix}0\\ 1/4\end{pmatrix}$ and form $\Omega$ as in Definition 3.3. Then $\Omega=\left\{\omega=\begin{pmatrix}\omega_{1}\\ \omega_{2}\end{pmatrix};-\frac{1}{2}<\omega_{1}<\frac{1}{2},0<\omega_{2}<\frac{1}{2}\right\}$ as illustrated in Figure 2. The open set $\sigma\Omega$ is outlined as well. Note that $\Pi\Omega=\Omega\cup\sigma\Omega$.
$\Omega$$\sigma\Omega$
Figure 2: The $\Gamma^{*}$-orbit of $\omega_{0}$ and domain $\Omega$.
Let’s construct a simple $\Pi$-invariant measurable range function for $\ell^{2}(\mathcal{L}^{*})$ based on $\Pi\Omega$. Let $\kappa=\begin{pmatrix}1\\ 0\end{pmatrix}\in\mathcal{L}^{*}$ and consider the one-dimensional subspace $V=\mathbb{C}(\delta_{0}+\delta_{\kappa})=\{\alpha(\delta_{0}+\delta_{\kappa}):\alpha\in\mathbb{C}\}$ of $\ell^{2}(\mathcal{L}^{*})$. Let $E$ be a measurable subset of $\Omega$ as pictured in Figure 3 where we have suppressed the vertical axis and changed the scale a little.
$E\sigma E$$E\!+\!\kappa$$\sigma(E\!+\!\kappa)$
Figure 3: The measurable subset $E$ of $\Omega$ and three of its shifts.
Now consider the map $J:\Pi\Omega\to Gr\big(\ell^{2}(\mathcal{L}^{*})\big)$ given by $J(\theta)=V$, for $\theta\in E\cup\sigma E$, and $J(\theta)=\{0\}$, for $\theta\in\Pi\Omega\setminus(E\cup\sigma E)$. Then, $\mathcal{M}_{J}$ consists of all $F\in L^{2}\big(\Pi\Omega,\ell^{2}(\mathcal{L}^{*})\big)$ such that $F(\theta)=0$, for a.e. $\theta\in\Pi\Omega\setminus(E\cup\sigma E)$, and $F(\theta)\in\mathbb{C}(\delta_{0}+\delta_{\kappa})$, for a.e. $\theta\in E\cup\sigma E$. Before looking at (10) in this situation, note that $x_{\rm id}=\begin{pmatrix}0\\ 0\end{pmatrix}$ and $x_{\sigma}=\begin{pmatrix}1/2\\ 0\end{pmatrix}$. So, $e^{2\pi i\nu\cdot x_{\rm id}}=1$ and $e^{2\pi i\nu\cdot x_{\sigma}}=(-1)^{\nu_{1}}$, for all $\nu=\begin{pmatrix}\nu_{1}\\ \nu_{2}\end{pmatrix}\in\mathcal{L}^{*}$. Consider the following three properties for an $f\in L^{2}(\mathbb{R}^{2})$:
(a) $\widehat{f}(\omega+\kappa)=\widehat{f}(\omega)$, for a.e. $\omega\in\Omega$.
(b) $\widehat{f}\big(\sigma(\omega+\kappa)\big)=-\widehat{f}\big(\sigma(\omega)\big)$, for a.e. $\omega\in\Omega$.
(c) $\widehat{f}(\theta)=0$, for all $\theta\in\mathbb{R}^{2}\setminus\big(E\cup(E+\kappa)\cup\sigma E\cup\sigma(E+\kappa)\big)$.
What Theorem 6.3 implies for the simple range function $J$ is that, if $\mathcal{V}$ denotes the set of all $f\in L^{2}(\mathbb{R}^{2})$ whose Fourier transform $\widehat{f}$ satisfies (a), (b), and (c), then $\mathcal{V}$ is a closed subspace of $L^{2}(\mathbb{R}^{2})$ that is invariant under shifts from the wallpaper group $pg$.
Although the range function used in Example 6.4 is simple compared to the complexity that is possible, the example does suggest that the statement of Theorem 6.3 can be refined. We return to $\Gamma$ denoting a crystal group in dimension $n$ with $\Pi$, $\mathcal{L}$, and $\Omega$ as before. For each $L\in\Pi$, define a unitary operator $U_{L}$ on $\ell^{2}(\mathcal{L}^{*})$ by $U_{L}h(\nu)=e^{2\pi i\nu\cdot x_{L}}h(\nu)$, for all $\nu\in\mathcal{L}^{*}$ and for each $h\in\ell^{2}(\mathcal{L}^{*})$. In Theorem 6.3, we used range functions based on $\Pi\Omega$ that are $\Pi$-invariant, so completely determined by their restriction to $\Omega$. We now give a version where $\Pi$-invariance is replaced with twisting by the unitaries $U_{L}$, $L\in\Pi$.
Definition 6.5.
For a measurable range function $J:\Omega\to Gr\big(\ell^{2}(\mathcal{L}^{*})\big)$, let $J^{\Gamma}:\Pi\Omega\to Gr\big(\ell^{2}(\mathcal{L}^{*})\big)$ be given by $J^{\Gamma}(L\omega)=U_{L}J(\omega)$, for each $\omega\in\Omega$ and all $L\in\Pi$.
If $J$ is a measurable range function for $\ell^{2}(\mathcal{L}^{*})$ based on $\Omega$, then it is clear that $J^{\Gamma}$ is a measurable range function for $\ell^{2}(\mathcal{L}^{*})$ based on $\Pi\Omega$. Note that we can create a $\Pi$-invariant range function $J^{\prime}$ from $J$ based on $\Pi\Omega$ by simply letting $J^{\prime}(L\omega)=J(\omega)$, for all $L\in\Pi$, $\omega\in\Omega$. We call $J^{\prime}$ the $\Pi$-invariant extension of $J$. The following lemma is immediate from the definitions.
Lemma 6.6.
Let $J$ be a measurable range function for $\ell^{2}(\mathcal{L}^{*})$ based on $\Omega$ and let $J^{\prime}$ be the $\Pi$-invariant extension. For any $G\in L^{2}\big(\Pi\Omega,\ell^{2}(\mathcal{L}^{*})\big)$, we have $G(L\omega)\in J^{\prime}(L\omega)$ if and only if $U_{L}\big(G(L\omega)\big)\in J^{\Gamma}(L\omega)$, for all $\omega\in\Omega$ and $L\in\Pi$.
We can now present a modified statement of Theorem 6.3 as a corollary.
Corollary 6.7.
Let $\mathcal{V}$ be a closed subspace of $L^{2}(\mathbb{R}^{n})$. Then $\mathcal{V}$ is invariant under shifts by elements of $\Gamma$ if and only if there exists a measurable range function $J$ for $\ell^{2}(\mathcal{L}^{*})$ based on $\Omega$ such that, for any $f\in\mathcal{V}$, there exists $F\in\mathcal{M}_{J^{\Gamma}}$ so that $\widehat{f}\big(L(\omega+\nu)\big)=\langle F(L\omega),\delta_{\nu}\rangle$, for almost every $\omega\in\Omega$, every $L\in\Pi$, and every $\nu\in\mathcal{L}^{*}$.
Proof.
Suppose that $\mathcal{V}$ is invariant under shifts by $\Gamma$, and let $K$ be a $\Pi$-invariant range function for $\ell^{2}(\mathcal{L}^{*})$ based on $\Pi\Omega$, as described in Theorem 6.3. Thus for any $f\in\mathcal{V}$, there exists $G\in\mathcal{M}_{K}$ such that $\widehat{f}\big(L(\omega+\nu)\big)=e^{2\pi i\nu\cdot x_{L}}\langle G(L\omega),\delta_{\nu}\rangle$ for a.e. $\omega\in\Omega$ and all $L\in\Pi$ and $\nu\in\mathcal{L}^{*}$. Let $J$ denote the restriction of $K$ to $\Omega$. Then $J^{\prime}=K$. Define $F\in L^{2}\big(\Pi\Omega,\ell^{2}(\mathcal{L}^{*})\big)$ by $F(L\omega)=U_{L}\big(G(L\omega)\big)$, for a.e. $\omega\in\Omega$ and any $L\in\Pi$. By Lemma 6.6, $F\in\mathcal{M}_{J^{\Gamma}}$. And
$\widehat{f}\big(L(\omega+\nu)\big)=e^{2\pi i\nu\cdot x_{L}}\langle G(L\omega),\delta_{\nu}\rangle=\langle U_{L}\big(G(L\omega)\big),\delta_{\nu}\rangle=\langle F(L\omega),\delta_{\nu}\rangle,$ (13)
for a.e. $\omega\in\Omega$, any $L\in\Pi$, and any $\nu\in\mathcal{L}^{*}$.
Conversely, suppose that for any $f\in\mathcal{V}$, there exists $F\in\mathcal{M}_{J^{\Gamma}}$ such that
$\widehat{f}\big(L(\omega+\nu)\big)=\langle F(L\omega),\delta_{\nu}\rangle$
for a.e. $\omega$ and all $L\in\Pi$, $\nu\in\mathcal{L}^{*}$. Defining $G\in L^{2}\big(\Pi\Omega,\ell^{2}(\mathcal{L}^{*})\big)$ by $G(L\omega)=U_{L}^{-1}F(L\omega)$ for a.e. $\omega\in\Omega$, it is immediate that $\langle F(L\omega),\delta_{\nu}\rangle=e^{2\pi i\nu\cdot x_{L}}\langle G(L\omega),\delta_{\nu}\rangle$, so that
$\widehat{f}\big(L(\omega+\nu)\big)=e^{2\pi i\nu\cdot x_{L}}\langle G(L\omega),\delta_{\nu}\rangle.$
We show that $G\in\mathcal{M}_{K}$, where $K$ is a $\Pi$-invariant measurable range function for $\ell^{2}(\mathcal{L}^{*})$ based on $\Pi\Omega$. Since $F\in\mathcal{M}_{J^{\Gamma}}$ by assumption, we have
$U_{L}G(L\omega)=F(L\omega)\in J^{\Gamma}(L\omega),$
which by Lemma 6.6 is equivalent to $G(L\omega)\in J^{\prime}(L\omega)$ for a.e. $\omega$, where $J^{\prime}$ is the $\Pi$-invariant extension of $J$. Thus $G\in\mathcal{M}_{J^{\prime}}$. It now follows from Theorem 6.3 that $\mathcal{V}$ is invariant under shifts by $\Gamma$. ∎